# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 30 Mar 2026 v3

@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Corollary 4 (Small-density moduli).
For every fixed $\tau\in(0,1/2)$, both sets $\{p:c(p)\leq\tau\}$ and $\{p:c(p)>\tau\}$ have positive relative density among the primes, equal to $G(\tau)$ and $1-G(\tau)$, respectively.
Proof.
By Theorem 7 with $\alpha=\tau$, the limiting density of $\{p:c(p)\leq\tau\}$ is $G(\tau)$. Since $G$ is strictly increasing on $0,\tfrac{1}{2}]$ by Theorem 6, we have $0<G(\tau)<1$, giving positive density for both sets. ∎
@@
Write
$c_{n}(p)=\frac{\varphi(p-1)}{p-1}\,P_{n}(p),\qquad P_{n}(p):=\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Since $P_{n}(p)\to 1$ as $p\to\infty$ through the primes, Lemma 5 shows that $c_{n}(p)$ has the same limiting law as $\varphi(p-1)/(p-1)$.
Since $c_{n}(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. Let
$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
For each odd prime $p\leq x$, we have
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
so $\omega(p-1)\leq m(x)$. Let
$r_{1}<r_{2}<\cdots<r_{d}$
be the distinct prime divisors of p1p-1, where $d=\omega(p-1)$. Then $d\leq m(x)$, and $r_{i}\geq p_{i}$ for $1\leq i\leq d$. Hence
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{d}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{d}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
because each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$. Also, by the same elementary estimate used in Lemma 1,
$P_{n}(p)\geq 1-\sum_{j=1}^{n}\frac{1}{p^{j}}\geq 1-\frac{1}{p-1}\geq\frac{1}{2}.$
Therefore
$c_{n}(p)\geq\frac{1}{2}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Now $N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}p_{m(x)+1}$, so $\log x\sim\log N_{m(x)}$. Since
$\log N_{m(x)}=\sum_{i=1}^{m(x)}\log p_{i}\sim p_{m(x)},$
we get $p_{m(x)}\sim\log x$. By Mertens’ theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x},$
and therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\gg\frac{1}{\log\log x}.$
For the upper bound, the argument of Proposition 1 applies verbatim, with the additional observation that $P_{n}(p)\leq 1$ for every prime $p$. If $p^{*}(k)\leq x$ is the least prime congruent to $1\pmod{N_{k}}$ arising there, then
$c_{n}\bigl(p^{*}(k)\bigr)=\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\,P_{n}\bigl(p^{*}(k)\bigr)\leq\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\ll\frac{1}{\log\log x}.$
Hence
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\ll\frac{1}{\log\log x},$
which proves the claim. ∎
5.3 Hausdorff dimension, moments, and Fourier structure of $G$
We derive a sharper geometric description of the limiting distribution and then its moment and Fourier structure. Throughout, $f$ is the strongly additive function defined in (1), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$. By Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
@@ 5.3 Hausdorff dimension, moments, and Fourier structure of $G$
By Tenenbaum [3, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
@@
For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$:
@@
If $\omega=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\omega\bigr)=e^{-I_{y}(\omega)},$
where $I_{y}(\omega)$ denotes the value of $I(y)$ at that configuration. Moreover,
@@
Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below). Next choose an increasing sequence $y_{n}\to\infty$ so rapidly that, with
$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2},$
the intervals $(y_{n},z_{n}]$ are pairwise disjoint, the sequence $(z_{n})$ is strictly increasing, and
$\sum_{n\geq 1}r_{n}<\infty,\qquad\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
@@
Let $T_{n}$ be the set of prefix configurations $\omega=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\omega)\leq Cn\log y_{n}$. Then
@@
Mertens’ theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
@@
Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$, and the second Borel–Cantelli lemma gives
@@
For each $\omega=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
$c(\omega):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
@@
$U_{n}:=\bigcup_{\omega\in T_{n}}[c(\omega),\,c(\omega)+r_{n}].$
@@
Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$. Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0. Therefore $\dim_{H}(\mu_{G})=0$. ∎
@@
(i) Set $k=1$ in (5): $\frac{\ell-2}{\ell-1}+\frac{1}{\ell}=\frac{\ell(\ell-2)+(\ell-1)}{\ell(\ell-1)}=\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}$.
(ii) Set $k=2$: $\frac{\ell-2}{\ell-1}+\frac{(\ell-1)}{\ell^{2}}=\frac{(\ell-2)\ell^{2}+(\ell-1)^{2}}{\ell^{2}(\ell-1)}=\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}$. (Verification at $\ell=3$: $\frac{13}{18}=\frac{1}{2}+\frac{2}{9}$. ) ∎
@@
It remains to verify analyticity and non-vanishing. Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$. The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\{\operatorname{Re}(s)>0\}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$ for all large $\ell$. Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g. [18, Ch. VII]) gives that the product converges uniformly on compact subsets to an analytic function. As $2^{-s}$ is entire, $M(s)$ is analytic on $\{\operatorname{Re}(s)>0\}$.
@@
$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$ (6)
@@ 5.4 Endpoint asymptotics at $1/2$
$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr).$
@@
6 Explicit Lower Bounds for Primes of Cryptographic Interest
We now give explicit lower bounds on $c(q)$ for primes $q$ whose predecessor q1q-1 has few distinct prime factors—a condition satisfied by many primes commonly used in lattice-based cryptography, especially Number Theoretic Transform (NTT)-friendly ones.
Theorem 11 (Explicit lower bound on $c(q)$ via the number of distinct prime factors).
Let $q\geq 3$ be a prime, and let $\omega=\omega(q-1)$ denote the number of distinct prime factors of q1q-1. Then
$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
@@ 6 Explicit Lower Bounds for Primes of Cryptographic Interest
Write the distinct prime factors of q1q-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
$\frac{\varphi(q-1)}{q-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
since the product is minimized by choosing the smallest possible distinct primes.
Now define
$P(t):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{t^{j}}\Bigr)\qquad(t>1).$
Each factor $1-t^{-j}$ is increasing in $t$, so $P(t)$ is increasing on $(1,\infty)$. Since $q\geq 3$ is prime, we have $P(q)\geq P(3)=P_{3}$. Therefore
$c(q)=\frac{\varphi(q-1)}{q-1}\,P(q)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
@@
The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 5: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$. Theorem 11 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order. The matching upper bound is Proposition 1. ∎
@@
Then for every prime $q\geq 3$,
$c(q)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $q$,
$c(q)\geq\frac{c_{0}}{\log\log q}.$
@@
$c(q)\geq P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
@@
$N_{\omega(q-1)}\leq\prod_{r\mid(q-1)}r\leq q-1,$
so $\omega(q-1)\leq m(q-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
$\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
@@
Finally, by Mertens’ theorem and Lemma 2,
$\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(q-1)}}\asymp\frac{1}{\log\log q},$
as $q\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(q)\ll\log\log q$. ∎
@@
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [19], Dusart [20]) with an explicit Linnik-type bound (see Xylouris [21]). We do not carry out the explicit computation here.
Remark 2.
The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large NTT primes with $q>2^{30}$, this factor differs from 11 by less than $10^{-9}$. Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [22, 23], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli. Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$. By Theorem 11, every prime $q\geq 3$ satisfies $c(q)\geq B(\omega(q-1))$. Direct evaluation gives the following values:
$\omega(q-1)\leq K$ $B(K)$
@@
Remark 3 (On the case $\omega(q-1)=1$).
If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$. Equivalently, $q$ is a Fermat prime. In this case
$\frac{\varphi(q-1)}{q-1}=\frac{1}{2},\qquad c(q)=\frac{1}{2}\,P(q),\qquad P(q):=\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr).$
Hence $c(q)$ is close to $1/2$ for large Fermat primes; so if infinitely many Fermat primes exist, then $c(q)\to 1/2$ along that subsequence. Thus the universal bound $B(1)=P_{3}/2\approx 0.280063$ is still not tight in this special case.
Corollary 10 (Practical bounds for cryptographic primes).
Let $q$ be a prime with $q>2^{30}$. Then
$c(q)>\frac{0.56}{\log q}.$
Moreover, if $\omega(q-1)\leq 10$, then
$c(q)>0.088.$
In particular, this applies to cryptographic-size primes $q\in[2^{64},2^{128}]$ satisfying $\omega(q-1)\leq 10$.
@@
The claim under $\omega(q-1)\leq 10$ follows from the bound $c(q)\geq B(\omega(q-1))$ (Theorem 11), since

### 21 Apr 2026 v4

@@ 5.3 Pure singularity
Write $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})$. Since $P_{n}(p)\to 1$, Lemma 5 gives the distributional claim. For the lower bound on $\min_{p\leq x}c_{n}(p)$: every odd prime $p\leq x$ satisfies $\omega(p-1)\leq m(x)$, so
$\frac{\varphi(p-1)}{p-1}\;\geq\;\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
and $P_{n}(p)\geq 1/2$ by Lemma 1. By Lemma 2 and Mertens’ third theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
The upper bound is identical to Proposition 1, since $P_{n}(p)\leq 1$. ∎
6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure? Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension. We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support is the full interval $0,\tfrac{1}{2}]$; in particular, the fractal-like clustering visible in Figure 2 is not an artifact of finite sampling. We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform. The section concludes with a precise asymptotic for the behavior of $G$ near the endpoint $1/2$.
Throughout, $f$ is the strongly additive function defined in (2), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$. By Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
@@ 6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
By Tenenbaum [38, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
@@
For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [17, Chapter 2]):
@@
If $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\boldsymbol{b}\bigr)=e^{-I_{y}(\boldsymbol{b})},$
where $I_{y}(\boldsymbol{b})$ denotes the value of $I(y)$ at that configuration. Moreover,
@@
Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below).
$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2}.$
We construct the sequence $(y_{n})$ recursively. Set $y_{1}:=3$. Given $y_{n}$, define $a_{n}$, $z_{n}$, $r_{n}$ as above and put $y_{n+1}:=z_{n}+1$. Then the intervals $(y_{n},z_{n}]$ are pairwise disjoint by construction, $z_{n}=y_{n}^{a_{n}}$ grows super-exponentially, and the summability conditions $\sum r_{n}<\infty$ and $\sum y_{n}^{-Cn}<\infty$ are immediate since $r_{n}=z_{n}^{-1/2}$ and $y_{n}^{-Cn}\leq y_{n}^{-C}$ with $y_{n}\to\infty$ super-exponentially.
@@
Let $T_{n}$ be the set of prefix configurations $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\boldsymbol{b})\leq Cn\log y_{n}$. Then
@@
Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
@@
Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$. Since the random variables $B_{\ell}$ are mutually independent (Lemma 10) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent. The second Borel–Cantelli lemma therefore gives
@@
For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
$v(\boldsymbol{b}):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
@@
$U_{n}:=\bigcup_{\boldsymbol{b}\in T_{n}}[v(\boldsymbol{b}),\,v(\boldsymbol{b})+r_{n}].$
@@
Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$. Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [17, Corollary 2.4]). Therefore $\dim_{H}(\mu_{G})=0$. ∎
The vanishing Hausdorff dimension of the limiting measure explains the clustering visible in Figure 2: the purely singular law forces the values of $c(p)$ to concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
Figure 2: Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Proposition 8 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$. The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$. The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 2.
@@
Immediate from (5) with $k=1$ and $k=2$. ∎
@@
It remains to verify analyticity and non-vanishing. Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$. The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\{\operatorname{Re}(s)>0\}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$ for all large $\ell$. Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g. [8, Ch. VII]) gives that the product converges uniformly on compact subsets to an analytic function. As $2^{-s}$ is entire, $M(s)$ is analytic on $\{\operatorname{Re}(s)>0\}$.
@@
$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$
@@
Figure 3: Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 5 using all primes $\ell\leq 10^{7}$. The top panel uses a linear scale and the bottom panel a log-log scale. The function oscillates but its local averages decrease across successive decades of $\tau$.
Remark 2 (Numerical evidence on Fourier decay).
We evaluated $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 5 with all primes $\ell\leq 10^{7}$. Figure 3 displays the result for $\tau\in[1,10^{6}]$. The function oscillates wildly, with local maxima near $0.14$ persisting at moderately large $\tau$. However, the mean value within each decade decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, $0.03$ for $10^{2},10^{3}]$, $0.02$ for $10^{3},10^{4}]$, $0.01$ for $10^{4},10^{5}]$, and $0.008$ for $10^{5},10^{6}]$. A power-law fit to the envelope of local maxima yields $|\widehat{\mu_{f}}(\tau)|^{2}\approx\tau^{-0.19}$. This is consistent with decay to 0, which would mean $\mu_{f}$ is a Rajchman measure; establishing this rigorously remains open (Section 8).
@@ 6.1 Endpoint asymptotics at $1/2$
The value $1/2$ is the supremum of the support of $G$ but is never attained (Remark 1). The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the rejection-sampling overhead $1/c(p)$ is close to its minimum value of $2$. We now determine this rate precisely.
@@
$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr)\approx 0.7413.$
@@
Remark 3 (Twin prime constant).
The constant $\kappa$ equals $C_{2}\,e^{-\gamma}$, where $C_{2}=2\prod_{p\geq 3}(1-(p-1)^{-2})\approx 1.3203$ is the twin prime constant of Hardy and Littlewood [22] (denote $2C_{2}$ therein). This follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $C_{2}/2$.
@@
7 Explicit Lower Bounds via the Number of Distinct Prime Factors
Theorem 9 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).
Let $p\geq 3$ be a prime, and let $\omega=\omega(p-1)$ denote the number of distinct prime factors of p1p-1. Then
$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
@@ 7 Explicit Lower Bounds via the Number of Distinct Prime Factors
Write the distinct prime factors of p1p-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$, we have $P(p)\geq P(3)=P_{3}$. Therefore, by (1),
$c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
@@
The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 4: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ third theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$. Theorem 9 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order. The matching upper bound is Proposition 1. ∎
@@
Then for every prime $p\geq 3$,
$c(p)\;\geq\;P_{3}\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $p$,
$c(p)\geq\frac{c_{0}}{\log\log p}.$
@@
$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
@@
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1,$
so $\omega(p-1)\leq m(p-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
$\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
@@
Finally, by Mertens’ third theorem and Lemma 2,
$\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(p-1)}}\asymp\frac{1}{\log\log p},$
as $p\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(p)\ll\log\log p$. ∎
@@
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [34, Theorems 4 and 7], Dusart [12, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [41, Theorem 1.1].
Remark 5 (Sharper bounds for specific primes).
Direct computation of the Euler product gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$. For the NIST moduli $p=3329$ and $p=8380417$ [31, 32], $\prod_{j\geq 1}(1-p^{-j})>0.9996$ and >0.9999998>0.9999998, respectively.
Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 9, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
$\omega(p-1)\leq K$ $B(K)$
@@
Table 1: Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 9. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
Remark 6 (On the case $\omega(p-1)=1$).
If $\omega(p-1)=1$ and $p$ is an odd prime, then p1p-1 is an even prime power, hence $p-1=2^{m}$ for some $m\geq 1$. Equivalently, $p$ is a Fermat prime. In this case
$\frac{\varphi(p-1)}{p-1}=\frac{1}{2},\qquad c(p)=\frac{1}{2}\,P(p),\qquad P(p):=\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Hence $c(p)\to 1/2$ along any infinite subsequence of Fermat primes, so $B(1)$ is not tight for Fermat primes.
Corollary 9 (Explicit lower bounds for large primes).
Let $p$ be a prime with $p>2^{30}$. Then
$c(p)>\frac{0.56}{\log p}.$
Moreover, if $\omega(p-1)\leq 10$, then
$c(p)>0.088.$
@@
The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 9), since